# From Pixels to Tokens: Byte-Pair Encoding on Quantized Visual Modalities

### v2

@@ Abstract
Multimodal Large Language Models have made significant strides in integrating visual and textual information, yet they often struggle with effectively aligning these modalities. We introduce a novel image tokenizer that bridges this gap by applying the principle of Byte-Pair Encoding (BPE) to visual data. Unlike conventional approaches that rely on separate visual encoders, our method directly incorporates structural prior information into image tokens, mirroring the successful tokenization strategies used in text-only Large Language Models. This innovative approach enables Transformer models to more effectively learn and reason across modalities. Through theoretical analysis and extensive experiments, we demonstrate that our BPE Image Tokenizer significantly enhances MLLMs’ multimodal understanding capabilities, even with limited training data. Our method not only improves performance across various benchmarks but also shows promising scalability, potentially paving the way for more efficient and capable multimodal foundation models.
@@ 1 Introduction
  • We design an algorithm for training the BPE Image Tokenizer and train an MLLM with this tokenizer. The performance evaluation further validates the capability enhancements this learning paradigm brings to MLLMs.
@@ 3 Theoretical Analysis
Here we provide a sketch of the proof of Proposition 1. For the full proof, please see Appendix A.
@@
The implications of Proposition 1 and Remark 1 are profound for understanding the limitations of simple models in capturing the structure of two-dimensional image data. This result shows a significant performance gap between unigram models and the optimal unconstrained model for such data. This gap, represented by the difference between $H(\pi)$ and $H_{\infty}$, can be arbitrarily large. This finding underscores a critical limitation of simple unigram models, which Transformers could default to when processing raw sequences, according to our earlier experimental results. These models are inherently restricted in their ability to capture the true structure of image data, suggesting that more sophisticated tokenization methods are necessary for effectively processing two-dimensional image data with Transformer models.
@@
Here we provide a sketch of the proof of Proposition 2. For the full proof, please see Appendix A.
@@
Proposition 2 offers valuable insights into the potential benefits of using a tokenizer for image data. This result demonstrates that with an appropriate tokenizer, we can achieve a loss arbitrarily close to the optimal unconstrained loss $H_{\infty}$, significantly outperforming the unigram model bound of $H(\pi)$. These theoretical results motivate the design of our BPE Image Tokenizer. In the following sections, we will describe the implementation of our algorithm and present corresponding experimental results.
@@ 4.1 BPE Image Tokenizer Training
Tokenizer Training Algorithm: We developed the BPE image tokenizer training algorithm, inspired by the text BPE algorithm. The main procedure of this algorithm is presented in Algorithm 4.1, with supporting functions detailed in Algorithms B.1 and B.2 (see Appendix B). Key designs of our algorithm include:
  • Use of an adjacency matrix to count token co-occurrences.
  • Simultaneous consideration of horizontal and vertical token combinations.
  • Ability to merge tokens into arbitrary shapes through iterative merging.
@@ 5.2 Impact of the BPE Image Tokenizer
Performance Gains: The integration of our BPE Image Tokenizer improves model performance across all benchmarks, as evidenced by comparing the ‘LLM+VQ’ and ‘LLM+VQ+BPE’ rows in Table 5.1. These gains indicate that the BPE Image Tokenizer provides the model with better multimodal understanding. The improvements is consistent across different training methods, also highlighting the robustness of our approach.
@@ 7 Conclusions, Limitations and Future Work
In this paper, we proposed a novel paradigm for MLLM training, demonstrating improvements in image understanding capabilities across various benchmarks, even with limited training data. Our approach centers on the novel BPE Image Tokenizer, which effectively combines image token IDs to facilitate better integration of visual information into MLLMs. We provided theoretical insights into the importance of tokenization for learning 2D sequence data with Transformer models, supporting our empirical findings and offering a new perspective on processing visual information in language models. Our experiments not only showcased the effectiveness of the BPE Image Tokenizer but also demonstrated the scalability of our approach.
@@
Appendix A Full Proofs
@@ B.4 Hyperparameters for training MLLM
Appendix C Examples
@@
Table C.1: Output examples of models trained with and without using BPE tokenizer. The text marked in red has obvious hallucinations.
@@
As shown in Table C.1, in these two cases, the version without using the BPE tokenizer, which directly uses VQ-quantized tokens for Llama-3.1-8B training, frequently exhibits hallucination issues. In contrast, the model trained with the BPE tokenizer can more accurately describe the information in the images, showcasing better image comprehension capabilities.
Appendix D Broader Impact
The development of our proposed noval paradigm for training MLLMs has potential to advance artificial intelligence and its applications across various domains, including medical imaging and autonomous systems. However, this progress also raises important ethical considerations, such as privacy concerns and the potential for bias in AI systems. As this technology evolves, it is crucial to balance innovation with responsible development, necessitating collaboration between researchers, ethicists, and policymakers. Future work should not only focus on technical improvements but also on ensuring that these advancements benefit society while mitigating potential risks. This research thus opens new avenues for AI capabilities while underscoring the importance of ethical considerations in technological progress.
Appendix E Computing Resources
We list the hardware resources used in Table E.1.
Table E.1: Computing resources for our experiments.

### v3

@@ Abstract
Multimodal Large Language Models have made significant strides in integrating visual and textual information, yet they often struggle with effectively aligning these modalities. We introduce a novel image tokenizer that bridges this gap by applying the principle of Byte-Pair Encoding (BPE) to visual data. Unlike conventional approaches that rely on separate visual encoders, our method directly incorporates structural prior information into image tokens, mirroring the successful tokenization strategies used in text-only Large Language Models. This innovative approach enables Transformer models to more effectively learn and reason across modalities. Through theoretical analysis and extensive experiments, we demonstrate that our BPE Image Tokenizer significantly enhances MLLMs’ multimodal understanding capabilities, even with limited training data. Leveraging this method, we develop Being-VL-0, a model that demonstrates superior performance across various benchmarks and shows promising scalability, potentially paving the way for more efficient and capable multimodal foundation models.
@@ 1 Introduction
  • We design an algorithm for training the BPE Image Tokenizer and train the Being-VL-0 with this tokenizer. The performance evaluation further validates the capability enhancements this learning paradigm brings to MLLMs.
@@ 3 Theoretical Analysis
Here we provide a sketch of the proof of Proposition 1. For the full proof, please see Appendix A.2.
@@
The implications of Proposition 1 and Remark 1 are profound for understanding the limitations of simple models in capturing the structure of two-dimensional image data. This result shows a significant performance gap between unigram models and the optimal unconstrained model for such data. This gap, represented by the difference between $H(\pi)$ and $H_{\infty}$, can be arbitrarily large. This finding underscores a critical limitation of simple unigram models, which Transformers could default to when processing raw sequences, according to our earlier experimental results. These models are inherently restricted in their ability to capture the true structure of image data, suggesting that more sophisticated tokenization methods are necessary for effectively processing two-dimensional image data with Transformer models. We further discuss the potential information loss of our tokenizer in Appendix A.1.
@@
Here we provide a sketch of the proof of Proposition 2. For the full proof, please see Appendix A.2.
@@
Proposition 2 offers insights into the potential benefits of using a tokenizer for image data. This result demonstrates that with an appropriate tokenizer, we can achieve a loss arbitrarily close to the optimal unconstrained loss $H_{\infty}$, significantly outperforming the unigram model bound of $H(\pi)$. These theoretical results motivate the design of our BPE Image Tokenizer. In the following sections, we will describe the implementation of our algorithm and present corresponding experimental results.
@@ 4.1 BPE Image Tokenizer Training
Tokenizer Training Algorithm: We developed the BPE image tokenizer training algorithm, inspired by the text BPE algorithm. The main procedure of this algorithm is presented in Algorithm 4.1, with supporting functions detailed in Algorithms B.1 and B.2 (see Appendix B). Key designs of our algorithm include: 1) Use of an adjacency matrix to count token co-occurrences; 2) Simultaneous consideration of horizontal and vertical token combinations; 3) Ability to merge tokens into arbitrary shapes through iterative merging.
@@ 5.2 Impact of the BPE Image Tokenizer
Performance Gains: The integration of our BPE Image Tokenizer improves model performance across all benchmarks, as evidenced by comparing the ‘LLM+VQ’ and ‘LLM+VQ+BPE’ rows in Table 5.1. These gains indicate that the BPE Image Tokenizer provides our Being-VL-0 model with better multimodal understanding. The improvements is consistent across different training methods, also highlighting the robustness of our approach.
@@ 7 Conclusions, Limitations and Future Work
In this paper, we proposed a novel paradigm for MLLM training, demonstrating improvements in image understanding capabilities across various benchmarks, even with limited training data. Our approach centers on the novel BPE Image Tokenizer, which effectively combines image token IDs to facilitate better integration of visual information into MLLMs. We provided theoretical insights into the importance of tokenization for learning 2D sequence data with Transformer models, supporting our empirical findings and offering a new perspective on processing visual information in language models. Our traind Being-VL-0 model not only showcased the effectiveness of the BPE Image Tokenizer but also demonstrated the scalability of our approach.
@@
Acknowledgments
This work was supported by NSFC under Grant 62450001 and 62476008. The authors would like to thank the anonymous reviewers for their valuable comments and advice.
Appendix A Theoretical Discussions
A.1 Information Loss of the tokenizer
Let $\bm{X}$ denote the original image, $\bm{V}$ denote the VQ-tokenized sequence, and $\bm{T}$ denote the final BPE-tokenized sequence. The information loss introduced by the BPE tokenization process can be quantified using conditional entropy:
$L_{bpe}=H(\bm{X}|\bm{T})-H(\bm{X}|\bm{V}).$ (6)
This quantity represents the increase in uncertainty about the original image $\bm{X}$ when using BPE tokens $\bm{T}$ compared to using VQ tokens $\bm{V}$. Since
$L_{bpe}=H(\bm{X}|\bm{T})-H(\bm{X}|\bm{V})=[H(\bm{X},\bm{T})-H(\bm{T})]-[H(\bm{X},\bm{V})-H(\bm{V})],$ (7)
and recall that mutual information can be expressed as:
$I(\bm{X};\bm{V})$ $=H(\bm{X})+H(\bm{V})-H(\bm{X},\bm{V}),$ (8)
$I(\bm{X};\bm{T})$ $=H(\bm{X})+H(\bm{T})-H(\bm{X},\bm{T}).$
Substituting back into equation 7 and we get
$L_{bpe}$ $=[H(\bm{X})+H(\bm{T})-I(\bm{X};\bm{T})-H(\bm{T})]-[H(\bm{X})+H(\bm{V})-I(\bm{X};\bm{V})-H(\bm{V})]$ (9)
$=[H(\bm{X})-I(\bm{X};\bm{T})]-[H(\bm{X})-I(\bm{X};\bm{V})]$
$=I(\bm{X};\bm{V})-I(\bm{X};\bm{T}).$
Therefore, the information loss can be expressed as the reduction in mutual information between the original image and its tokenized representation. This formulation is particularly useful as it directly quantifies how much information about the original image is preserved through the tokenization process.
Next, we need to analyze how each BPE merge operation affects the mutual information $I(\bm{X};\bm{T})$. For any given BPE token sequence $\bm{T}$, we can decompose $H(\bm{X}|\bm{T})$ based on individual tokens:
$H(\bm{X}|\bm{T})=\sum_{t}P(t)H(\bm{X}|\bm{T}=t).$ (10)
During a merge operation that combines tokens $(t_{i},t_{j})$ into $t_{m}$, the change in conditional entropy can be expressed as:
$\Delta H=P(t_{m})H(\bm{X}|\bm{T}=t_{m})-[P(t_{i},t_{j})H(\bm{X}|\bm{T}=t_{i},t_{j})].$ (11)
For the whole token distribution, the increase in conditional entropy can be quantified using the KL divergence:
$H(\bm{X}|t_{m})=H(\bm{X}|t_{i},t_{j})+\mathrm{KL}(P(\bm{X}|t_{i},t_{j})||P(\bm{X}|t_{m})),$ (12)
where $\mathrm{KL}(\cdot)$ represents the KL divergence. In this equation,
  • $H(\bm{X}|t_{i},t_{j})$ represents the original uncertainty about $\bm{X}$ given the separate tokens
  • $\mathrm{KL}(P(\bm{X}|t_{i},t_{j})||P(\bm{X}|t_{m}))$ represents the extra uncertainty introduced by merging the tokens
  • The equation shows that no other sources of information loss exist beyond these two terms.
This leads to an initial upper bound on the total information loss:
$L_{bpe}\leq\sum_{merges}\mathrm{KL}(P(\bm{X}|t_{i},t_{j})||P(\bm{X}|t_{m})).$ (13)
The summation here represents the sum over all merges during the BPE process. Considering that BPE performs merges based on frequency patterns and with minimum merge frequency threshold $p_{\min}$, we can relax the inequality to get a more interpretable bound:
$L_{bpe}\leq(|D_{bpe}|-|D_{vq}|)\times(-p_{\min}\log(p_{\min})).$ (14)
Here, $|D_{vq}|$ is the size of the VQ codebook, and $|D_{bpe}|$ is the size of the vocabulary after BPE extension.
To put this bound in perspective, consider a typical configuration:
  • $|D_{vq}|=8192$ (VQ codebook size)
  • $|D_{bpe}|=8192+8192$ (vocabulary size after BPE extension)
  • $p_{\min}=0.01$ (minimum merge frequency)
The upper bound on information loss for the whole vocabulary would be:
$L_{bpe}\leq(8192+8192-8192)\times(-0.01\times\log(0.01))\approx 377.3~{}\mathrm{bits}.$ (15)
For a single image, the original VQ tokens ($32\times 32$ patches) contain $32\times 32\times\log_{2}8192=13312~{}\mathrm{bits}$ information, and the per token loss of the extended BPE vocabulary is $L_{bpe}/(|D_{bpe}|-|D_{vq}|)\approx 0.046~{}\mathrm{bits}$. We can calculate that the max loss ratio of the single image is only $32\times 32\times 0.046/13312\approx 0.35\%$, which is a relatively small information loss. Considering the benefits brought by BPE tokenization as discussed earlier, we believe this loss is acceptable.
A.2 Full Proofs
@@ B.4 Hyperparameters for training MLLM
B.5 Details of the Pipeline
In this section, we provide more detailed descriptions of our pipelines as below.
B.5.1 LLM+VQ+BPE Pipeline
Our complete pipeline (LLM+VQ+BPE) consists of three main components:
  • Image Quantization: We employ a pretrained VQ-GAN model (with a codebook size of 8192) to quantize input images into a sequence of token IDs. This process converts images of any size into a 1024-dimensional ID tensor.
  • BPE Image Tokenizer: This component, which distinguishes our approach from previous methods, processes the quantized token IDs using our trained BPE tokenizer. The tokenizer learns to merge frequently co-occurring tokens, creating a new vocabulary that captures higher-level visual patterns.
  • Multimodal Learning: The processed image token IDs are combined with text token IDs and fed into the LLM (Llama-3.1-8B in our experiments). Before training, we expand the token embedding layers of the base model to accommodate both the VQ-GAN token IDs (8192) and the BPE vocabulary (e.g., 4096), along with special tokens for marking image positions.
B.5.2 LLM+VQ Baseline
The LLM+VQ variant serves as our baseline, following a simpler approach:
  • It uses the same VQ-GAN model for image quantization as in LLM+VQ+BPE.
  • The quantized token IDs are directly combined with text token IDs, bypassing the BPE tokenizer stage.
  • The token embedding layers are only expanded to accommodate the VQ-GAN token IDs and special tokens.
B.5.3 Training Process
Both variants undergo the same two-stage training process:
  • Image Understanding Pretraining (PT): During this stage, we train the expanded token embeddings using image caption data. In the PT(freeze) setting, we freeze the original text token embeddings, while in PT(full), all embeddings are updated.
  • Supervised Fine-Tuning (SFT): All parameters are unfrozen and fine-tuned using SFT data.
Appendix C Additional Results
C.1 Additional Training with Llama2 version / freezing during SFT
To further investigate the effects of different base models and varying component freezing strategies, we conducted additional experiments. The results are shown in Table C.1
Table C.1: Full results with additional experiments. The red parts indicate using Llama2 as the base LLM for training. The blue parts indicate freezing the embedding corresponding to the visual token during the SFT stage.
Training type VQAv2 MMBench MMEp MMEc POPE VizWiz
Llama3.1+VQ SFT 51.1 35.9 972.3 231.8 73.8 43.1
PT(full)+SFT 53.7 37.0 1037.2 261.4 75.3 44.2
PT(freeze text)+SFT 55.4 37.6 1054.5 277.0 76.0 45.3
Llama2+VQ PT(freeze text)+SFT 54.0 35.7 991.9 254.2 75.1 44.4
Llama3.1+VQ+BPE SFT 52.2 35.4 1029.7 269.6 76.3 45.3
PT(full)+SFT(freeze visual) 31.7 17.8 624.1 171.9 46.4 29.5
PT(full)+SFT 56.5 38.6 1144.6 284.3 77.3 45.8
PT(freeze text)+SFT(freeze visual) 22.5 13.3 488.7 143.6 35.2 21.5
PT(freeze text)+SFT 57.1 40.9 1223.5 307.1 79.0 46.0
Llama2+VQ+BPE PT(freeze text)+SFT 56.5 38.1 1112.2 277.8 77.9 44.9
+RefCOCO(50.6K) 58.6 42.3 1257.4 314.3 79.8 47.1
+ShareGPT4o (57.3K) 60.2 43.7 1304.5 327.7 80.9 47.8
C.2 Additional Evaluation
To assess our method’s effectiveness across various capabilities, we performed comprehensive evaluations analyzing performance across distinct sub-categories. The results are shown in Table C.2 and C.3
Table C.2: Comparison of LLM+VQ+BPE and LLM+VQ performance across different categories in MME benchmark.
Category LLM+VQ+BPE LLM+VQ
Perception Existence 145.00 113.33
Count 120.00 110.00
Position 106.67 103.33
Color 148.33 120.00
Posters 136.24 121.08
Celebrity 111.76 89.51
Scene 125.00 101.75
Landmark 130.25 110.00
Artwork 112.75 90.50
OCR 87.50 95.00
Cognition Commonsense Reasoning 107.14 89.57
Numerical Calculation 62.50 50.00
Text Translation 95.00 102.50
Code Reasoning 42.50 35.00
Table C.3: Comparison of LLM+VQ+BPE and LLM+VQ performance across different categories in MLLM-bench. The numbers in the table represent the quantity of answers judged to be better for each respective model. “Tie” indicates the number of answers where there is no significant difference between the two models’ answers.
Category LLM+VQ+BPE LLM+VQ Tie
perception 26 34 10
understanding 52 33 25
Applying 27 14 19
Analyzing 49 40 31
Evaluation 12 19 9
Creation 7 9 4
Total 173 149 98
Table C.4: Evaluation results on the Nocaps and Flickr30k benchmarks. For Nocaps, we chose the validation split. For Flickr30k, we selected a 1k-image split. We compared the impact of adding BPE on the performance of image captioning tasks that require detailed understanding. Additionally, we investigated versions of LLM+VQ+BPE (+SFT), which used additional scaling as described earlier, to observe how more instruction tuning affects performance in image captioning tasks.
Model Nocaps Flickr30k
CIDEr SPICE METEOR ROUGE-L CIDEr SPICE METEOR ROUGE-L
LLM+VQ 76.4 17.2 24.4 51.3 76.4 16.2 24.4 51.3
LLM+VQ+BPE 75.7 16.9 24.1 50.8 75.7 15.9 24.1 50.8
LLM+VQ+BPE (+SFT) 80.7 18.2 25.0 52.6 80.7 17.2 25.0 52.6
Appendix D Additional Discussion about Further Scaling
Our methods with BPE Image Tokenizer represent a departure from conventional CLIP-based visual encoding methods, introducing both challenges and opportunities that warrant careful discussion.
A critical aspect of evaluating our approach is understanding the implicit data scale disparity between our method and existing CLIP-based MLLMs. Traditional visual encoders typically leverage massive pretraining datasets—CLIP’s original implementation utilized 400M image-text pairs (Radford et al., 2021), while subsequent models expanded to datasets like LAION-2B, processing billions of samples across multiple epochs. In contrast, our BPE Image Tokenizer achieves competitive performance using approximately 2.78M images, representing roughly 0.1% of the training data used by encoders like CLIP-ViT-L/14.
This substantial difference in pre-training data requirements highlights a fundamental trade-off in MLLM development: while CLIP-based approaches benefit from extensive pre-training, our method demonstrates remarkable efficiency in learning visual representations from limited data. The comparable performance achieved with significantly fewer resources suggests that our approach may offer a more scalable path forward for MLLM development, particularly in scenarios where massive datasets or computational resources are not available.
While our current implementation may not achieve state-of-the-art performance compared to heavily optimized CLIP-based MLLMs, this limitation should be viewed within the broader context of our research objectives. The primary contribution of this work lies not in establishing a new performance benchmark, but rather in introducing and validating a novel training paradigm for MLLMs. The theoretical foundations and experimental results presented here serve as proof-of-concept for an alternative approach to visual-language modeling.
We acknowledge that achieving state-of-the-art performance would require:
  • Substantial expansion of training data
  • Significant computational resources for large-scale training
  • Extensive engineering optimizations
These requirements extend beyond the scope of our current investigation, which focuses on establishing the theoretical and practical viability of our approach.
The promising results obtained with limited resources suggest exciting potential for future research. We believe our work lays the groundwork for a new direction in MLLM research, demonstrating that alternative approaches to visual-language modeling can achieve competitive performance even with limited resources. As the field continues to evolve, the principles and methods introduced in this paper may contribute to the development of more efficient and scalable MLLMs. Our future work will focus on scaling up these initial findings while preserving the core benefits of our approach.
Appendix E Examples
@@
Table E.1: Output examples of models trained with and without using BPE tokenizer. The text marked in red has obvious hallucinations.
@@
As shown in Table E.1, in these two cases, the version without using the BPE tokenizer, which directly uses VQ-quantized tokens for Llama-3.1-8B training, frequently exhibits hallucination issues. In contrast, the model trained with the BPE tokenizer can more accurately describe the information in the images, showcasing better image comprehension capabilities.
Appendix F Broader Impact
The development of our proposed novel paradigm for training MLLMs has the potential to advance artificial intelligence and its applications across various domains, including medical imaging and autonomous systems. However, this progress also raises important ethical considerations, such as privacy concerns and the potential for bias in AI systems. As this technology evolves, it is crucial to balance innovation with responsible development, necessitating collaboration between researchers, ethicists, and policymakers. Future work should not only focus on technical improvements but also on ensuring that these advancements benefit society while mitigating potential risks. This research thus opens new avenues for AI capabilities while underscoring the importance of ethical considerations in technological progress.
Appendix G Computing Resources
We list the hardware resources used in Table G.1.
Table G.1: Computing resources for our experiments.